# RealCQA-V2 : Visual Premise Proving thanks: WIP: Data will be uploaded soon

### v1

@@ Abstract
We introduce Visual Premise Proving (VPP), a novel task tailored to refine the process of chart question answering by deconstructing it into a series of logical premises. Each of these premises represents an essential step in comprehending a chart’s content and deriving logical conclusions, thereby providing a granular look at a model’s reasoning abilities. This approach represents a departure from conventional accuracy-based evaluation methods, emphasizing the model’s ability to sequentially validate each premise and ideally mimic human analytical processes. A model adept at reasoning is expected to demonstrate proficiency in both data retrieval and the structural understanding of charts, suggesting a synergy between these competencies. However, in our zero-shot study using the sophisticated MATCHA model on a scientific chart question answering dataset, an intriguing pattern emerged. The model showcased superior performance in chart reasoning (27%) over chart structure (19%) and data retrieval (14%). This performance gap suggests that models might more readily generalize reasoning capabilities across datasets, benefiting from consistent mathematical and linguistic semantics, even when challenged by changes in the visual domain that complicate structure comprehension and data retrieval. Furthermore, the efficacy of using accuracy of binary QA for evaluating chart reasoning comes into question if models can deduce correct answers without parsing chart data or structure. VPP highlights the importance of integrating reasoning with visual comprehension to enhance model performance in chart analysis, pushing for a balanced approach in evaluating visual data interpretation capabilities.
@@ 1 Introduction
Understanding and reasoning with visual data is crucial for Multi-Modal Question Answering (MMQA) systems. However, current methods do not effectively evaluate the sequential reasoning needed for complete comprehension, resulting in a gap between visual data representation and semantic interpretation. Current benchmarks in structured visual reasoning, such as those used in [1, 2], primarily rely on one-to-one accuracy metrics, which do not adequately capture the true reasoning capabilities of these models. While formal reasoning tasks like Conjecture/Theorem Proving and First-Order Logic (FOL) Verification have been extensively studied in structured domains like mathematics and natural language processing, such formalization is largely absent in vision-language MMQA settings.
@@
  • We propose two novel evaluation metrics, $ACC_{VPP}$ and $DCP$, to provide interpretable and explainable assessments of visual reasoning capabilities.
  • We curate a challenging dataset comprising 10 million real scientific chart question-answer pairs, annotated with premise-conclusion sequences, for evaluating sequential reasoning.
@@ 1.1 Problem Definition: Visual Reasoning Chains
The formalization of generalized visual reasoning is still in its early stages and involves tackling complex, higher-order reasoning problems. To make this problem tractable, we introduce two key constraints:
First, we limit the scope to First-Order Logic (FOL) problems. FOL verification is inherently complex and often NP-hard, particularly when dealing with natural language, which have inherent lexical, syntactic, and contextual ambiguities unless manually annotated.
Second, we focus specifically on Chart Question Answering (CQA). The recently released RealCQA dataset [3] serves as an ideal testbed for this task. This dataset, which is template-based (avoiding complex language), manually annotated (ensuring consistency and completeness), and grounded in mathematical logic (ensuring coherent quantification), makes the process of logical verification more manageable.
By restricting our problem to this structured and controlled environment, we establish a foundational framework for formalizing visual reasoning in MMQA. Our current work begins with chart QA and aims to extend to general visual question answering (VQA) and, eventually, more abstract visual reasoning in future works. The primary challenge lies in identifying and verifying visual premises and reasoning chains in diverse and unconstrained environments. Our work takes a step in this direction by focusing on the constrained domain of Chart Question Answering, where we introduce the VPP task to evaluate a model’s reasoning capacity in a manner analogous to human chart analysis. This not only measures the model’s final answer accuracy but also its ability to replicate the reasoning process across multiple steps.
@@ 2.1 Vision-Language Models
Figure 1: Diverse levels of visual reasoning are required for robust chart analysis.
@@ 2.2 Visual Reasoning over Charts
Charts convey complex information through visual elements like trend lines and legends, requiring detailed interpretation of various visual tasks (see Figure 1). These require prediction and evaluation at each step and are referred to as ‘Dense’ in Table1. Chart-to-text tasks like QA and summarizing involve directly evaluating chart understanding tasks through the text output.
Dataset # Img Source Task
FigureSeer [10] 1k ArXiv Dense
UB-PMC [11] 28k PubMed Central Dense
Real CQA [3] 28k PubMed Central QA
ChartQA [1] 22k Pew/Statista/OWID QA
C2T [12] 82k Pew/Statista Summary
EC400k [13] 400k Excel Dense
FQA [14] 180k Synthetic QA
DVQA [14] 300k Synthetic QA
LeafQA [15] 200k Synthetic QA
Table 1: Popular Chart Datasets, Sources and Tasks
@@
2.2.1 Chart Datasets
2.2.2 Chart Understanding Models :
Dense prediction models like Cached [16] focus on chart structure prediction and proposed a custom local-global context modelling architecture based on Swin-Transformers. Chart-Ocr[13] and SpaDen [17] focuses on using keypoints for tabular data extraction, by recreating plot area components like bars, lines etc and end-to-end data extraction with legend mapping.
Works focusing on chart question answering, like CRCT [18] use classification and regression predictors based on ‘Answer Type’. ChartQA[1] introduced template and human annotated question answers, baselined through offline tabular extraction using ChartOCR and answer prediction using VLT5. Matcha [19] utilizes the Pix2Struct model[6] with training on HTML structure, Math and ChartQA to get an impressive jump in $20\%$ accuracy on the ChartQA dataset.
To tackle Chart to Text tasks, Chartllama [20] first proposed using a generic visual encoder with LLAMA. ChartReader [21] further leverages a keypoint-based chart component detector with a T5-based text decoder. StructChart [22] showed novelty by reformulating chart data into triplets, effectively capturing both the structure and semantics inherent in the data. ChartAssistant [23] integrates Swin and ViT encoders with Bart and LLAMA language decoders, evaluating across multiple datasets. All of these chart to text works suffer from the inherent ambiguity in one-to-one evaluation of QA based visual reasoning.
@@ 3 Chart FOL Premise Proving
(a) Structure Premises
(b) True/False Count
(c) PCA
(d) tSNE
We discuss our methodology for the proposed task. First we discuss our base ideation and task formalization, then discuss some meta charecteristics of the created data from the approach and finally our proposed evaluation metric for the task.
We want to build a system for evaluation of visual reasoning from first principles as depicted in Figure 2. Our basic idea derives this theory from the way human beings read semantically arranged visual representations of data, i.e. charts. We start from the input an Image $I\in R^{3}$ and a question. Each question context is confined to the template variables and predicates. This can be systematically characterized into a series of premise-conclusions per question, visually grounded in the plot area.
@@ 3.1 VPP Task
"We start by defining a Premise-Conclusion Pair, where each premise logically leads to a conclusion. These premises are articulated in FOL, utilizing quantifiers such as ‘for all’ ($\forall$) and ‘there exists’ ($\exists$), alongside predicates applicable to these variables. Subsequently, we establish a Chain of Reasoning. This sequence involves linking each proposition to the next, ensuring that each step logically follows from the previous one as described in Algorithm 1 and the example in 3.2.1, and 3.2.2 . The task culminates in Proving the Premises, where the final premise-conclusion, the original question—answer is derived from the interconnected premises through the established reasoning chain. Successfully completing this task confirms that the conclusion is a logical consequence of the premises. Since each premise is an FOL statement we can represent our final premise conclusion as natural language strings, as well as mathematical graph representations [24] using abstract syntax trees (AST) uniquely identifying each logical sequence. Representing textual premises as graphs where nodes correspond to entities or concepts mentioned in the premises, and edges represent logical or causal relationships between these entities enables problem formulation as a graph problem for visual logic verification studied independantly of the natural language complexity. While we have released the AST representation of our dataset we currently focus on the NLP version for this work.
@@ 3.2 Premise Conclusion Sequence Creation
We define our step by step process of identifying and creating visual premises for charts in Algorithm 1. We create four types of chart premises:
@@
    Structural Premises (SP) Identify the chart’s structure, categorized into 10 types (e.g., 0: Chart Type, 1: Y-Title, 2: X-Title, 3: Y-Range, 4: X-Range, 5: Categorical, 6: X-Tick, 7: Y-Tick, 8: # of Dataseries, 9: Legend Label as shown in Fig 3(a)).
@@
    Data Premises (DP) per chart image for data retrieval verification of that plot area.
@@
    Reasoning Premises (RP) per question template requiring logical deduction based on chart data and explicit calculations like comparing or transforming data values.
@@
Thus each question template is deconstructed to sequential premise templates /AST. Each premise template is used to generate a premise string and has true or false conclusion specific to query values of that chart used to generate the string. For True conclusions we use actual variable value present in the groundtruth chart annotation. For False conclusions we randomly use any other relevant variable values present in the current chart (eg. out of range tick value, axis title instead of legend label etc). This setup effectively breaks down the original question into a comprehensible set of binary premise-conclusions, generated in a ratio of 1:3 true-to-false for each as shown in Figure 3(b). This is done by design to bias the dataset with distractor (False) representations and prevent overfitting to actual (True) conclusion statements. In generating the final Premise strings we further add vocabulary complexity by choosing at random 1 of 3 paraphrased templates created using a pre-trained T5-Transformer from original handcrafted premise template.
The PCA and t-SNE plots of pretrained-BERT embeddings as shown in Figures 3(c), and 3(d) for 25,364 unique words indicate a highly diverse vocabulary, with a broad and evenly distributed semantic space and no apparent clustering, highlighting rich semantic coverage. In contrast to the focused and less diverse vocabularies in popular VQA tasks, which show distinct clusters around common categories and actions, this corpus encompasses a wider range of topics and semantics. The extensive diversity and rich semantic distribution suggest a more nuanced and challenging dataset.
Algorithm 1 Parsing a Chart from First Principles
Input: QA template, FOL templates
Parameters: Chart structure variables per image ($x,y,\ell,t$), QA values per string ($v,d$), FOL predicates
Output: Sequence of FOL premise-conclusion pairs
1:  while Chart components {$x,y,\ell,t,v,d$} exist do
2:     Recognize axis labels and verify variables: $\forall x\in\text{Axis Labels},\exists y\in\text{Variables}:\text{InRange}(y)$
3:     if Legend exists then
4:        Locate legend and map data series: $\exists\ell\in\text{Legends}:\text{LocateRecognize}(\ell)$
5:     end if
6:     Verify tick marks on axis: $\forall t\in\text{Tick Marks},\text{Valid}(t)\wedge\neg\text{BrokenAxis}(t)$
7:     Verify data values on plot area: $\forall v\in\text{Tick Values},\exists d\in\text{Data Points}:\text{AtLocation}(v,d)$
8:     if Legend exists then
9:        Check data values for each legend: $\forall v\in\text{Tick Values},\ell\in\text{Legends},\exists d\in\text{Data Points}:\text{AtLocationWithLegend}(v,\ell,d)$
10:     end if
11:  end while
12:  Verify predicate logic for validated data values
13:  if Implicit math transformation required then
[…]
15:  end if
3.2.1 Example Query Template
Is the difference between the value of $Y$ for legendlabel1 at the $i$-th $x$ tick and the $j$-th $x$ tick greater than the difference between the value of $Y$ for legendlabel2 at the $i$-th $x$ tick and the $j$-th $x$ tick?
3.2.2 Deconstructed Sequential Premises
[…]

### v2

# RealCQA-V2 : Visual Premise Proving A Manual COT Dataset for Charts thanks: WIP: Data will be uploaded soon

@@ Abstract
Large vision-language models (LVLMs) have shown promise in multimodal reasoning, yet they lack a structured approach to formalize visual reasoning—a crucial gap for high-stakes tasks requiring logical rigor. To address this, we introduce RealCQA-V2, a dataset and task framework that bridges this gap by formalizing reasoning in chart-based question answering (Chart QA). Through Visual Premise Proving (VPP), we decompose reasoning into distinct logical premises, each capturing a necessary step for understanding and analyzing visual chart data. This approach extends the paradigm of Chain of Thought (COT) prompting into the visual domain, fostering interpretable and stepwise validation of reasoning processes.
RealCQA-V2 comprises over 10 million question-answer pairs derived from real-world scientific charts, annotated with premise-conclusion sequences to facilitate logical consistency in visual reasoning chains. We introduce two novel metrics—ACCVPP (Accuracy of Proven Premises) and DCP (Depth of Correct Premises)—to evaluate model performance on both correctness and reasoning depth, providing a holistic assessment beyond final-answer accuracy. Our experiments reveal that while LVLMs show potential in premise validation, they struggle with consistency in extended reasoning chains, highlighting the challenges and opportunities for improvement in logical coherence.
By formalizing visual reasoning in Chart QA, RealCQA-V2 paves the way for future research on interpretable and logically structured reasoning in large-scale vision-language models, advancing toward the goal of formal visual reasoning across multimodal AI.
@@ 1 Introduction
Understanding and reasoning with visual data is crucial for Multi-Modal Question Answering (MMQA) systems. However, current methods do not effectively evaluate the sequential reasoning needed for complete comprehension, resulting in a gap between visual data representation and semantic interpretation. Current benchmarks in structured visual reasoning, such as those used in [1, 2], primarily rely on one-to-one accuracy metrics, which do not adequately quantify the actual reasoning capabilities of these models. The era of large language models have marked a shift towards finding traction on reasoning based tasks. While showing impressive performance, their ability to navigate increasingly complex tasks have been further improved by leveraging intermediate results. This has been termed ‘Chain-of-Thought’ (COT) Prompting to generate a logical sequence of rationales for guidance to the final answer [3, 4, 5]. While formal reasoning tasks like Conjecture/Theorem Proving and First-Order Logic (FOL) Verification have been extensively studied in structured domains like mathematics and natural language processing, such formalization is largely absent in vision settings.
@@
  • We propose two novel evaluation metrics, $ACC_{VPP}$ and $DCP$, to provide interpretable and explainable assessments of COT reasoning capabilities.
  • We curate a challenging Manual-COT dataset comprising 10 million real scientific chart question-answer pairs, annotated with premise-conclusion sequences, for evaluating sequential reasoning.
@@ 1.1 Problem Definition: Visual Reasoning Chains
The formalization of generalized visual reasoning is still in its early stages and involves tackling complex, higher-order reasoning problems. To make our study tractable, we introduce two key constraints:
First, we limit the scope to First-Order Logic (FOL) problems. FOL verification is inherently complex and often NP-hard, particularly when dealing with natural language, which have inherent lexical, syntactic, and contextual ambiguities unless manually annotated. Second, we focus specifically on visual complexity of real world Chart Question Answering (CQA). The recently released RealCQA dataset [6] serves as an ideal testbed for this task. It is designed with a template-based approach to simplify language complexity, incorporates manually annotated visual components to ensure consistency and completeness, and is grounded in mathematical logic for coherent quantification. Together, these features make the process of logical verification in visual reasoning tasks more manageable.
By restricting our problem to this structured and controlled environment, we establish a foundational framework for formalizing visual reasoning in MMQA. Our current work begins with chart QA and aims to extend to general visual question answering (VQA) and, eventually, more abstract visual reasoning in future works. The primary challenge lies in identifying and verifying visual premises and reasoning chains in diverse and unconstrained environments. Our work takes a step in this direction by focusing on the constrained domain of Chart Question Answering, where we introduce the VPP task to evaluate a model’s reasoning capacity in a manner analogous to human chart analysis. This not only measures the model’s final answer accuracy but also its ability to replicate the reasoning process across multiple steps. The main motivations for our work are improving :
  • Multi-modal COT
    • Explainable Evaluation Metric
    • Formal Reasoning Framework
  • Real World Chart Reasoning
@@ 2.1 Vision-Language Models
Figure 1: Tasks required for complete chart analysis.
2.2 Large Models and COT
The initial research on Chain-of-Thought Reasoning, observed the phenomenon is exhibited prominently in larger models ($>100$bil params)[3]. Smaller models, while fluent in language, tended to produce illogical reasoning chains. Further on techniques like knowledge distillation[13, 14] and iterative prompting [15] have been shown to mitigate this degradation and elicit COT by fine-tuning as well. The NLP modality [16, 17, 18, 19] has been a popular focus. COT can be implemented through either crafting demos by human experts (Manual COT) [3], or through automated generation(Auto-COT) [5]. Manual-COT Datasets are very expensive and painstaking to curate, Auto-COT however suffers from hallucination. Demos are applied in zero-shot or few-shot settings.
Multimodal COT is a promising direction. To create multimodal LLM’s, techniques like transforming different modalities into a unified format and feeding them into LLMs [20, 21] or fine-tuning smaller language models with fused multimodal features [22, 23] have been explored. Recent works, [24, 25] introduced multimodal COT, that separates rationale generation and answer inference.
Multimodal COT Benchmarks have been proposed recently with annotated reasoning chains. A-OKVQA [26] consists of images from COCO-17 and crowd-sourced QA, annotators were tasked to make QA diverse. PMR[27] constructs premises by crowd-sourcing QA and taking images primarily of human interactions from movie screenshots, annotators are given premise templates. ScienceQA [28] consists of reasoning at a grade 1-12 level for school lectures, thus QA is structured academic knowledge. While these works have tried to capture a broad based knowledge setting required to reason, their evaluation is only through the accuracy of the final answer. As noted by researchers [24] there is speculation to the amount of hallucination in the rationale generation. Even though there is a plethora of research on COT to improve reasoning they fail to quantify the extent of correctness in a logical sequence only reporting aggregate final accuracy [28, 26]. Furthermore working in abstract higher-order reasoning space [27] makes it hard to formalize reasoning sequences, are the given premises the only steps or there exist other non-overlapping steps to reach same conclusion etc. ie How to measure the atmoicity or canonical logic path to a solution.
Formalization of COT remains a significant challenge towards meaningful quantification of reasoning capabilities, with the NLP community leading efforts in this area. One approach translates logical forms into natural language templates, constraining to counterfactual reasoning in real-world factual data [29]. Another explores reasoning in fictional contexts but lacks fine-grained analysis of intermediate COT steps [30]. Studies have shown that LLM’s struggle with planning tasks, although it is unclear if this stems from reasoning limitations or incomplete world modeling [31]. These challenges inspired PrOntoQA[32], which isolates reasoning evaluation by using fictional ontologies to verify each logical step while providing a formal analysis framework for reasoning ability over current COT datasets. It offers a method to assess validity, atomicity or canonical structure, and utility of each proof step, ensuring that models not only reach the correct conclusion but do so through relevant and logically sound steps. This framework’s reliance on synthetic data, however, limits its applicability to real-world datasets, which lack predefined logical paths. Without structured steps, verifying atomic or canonical steps in real-world data becomes challenging, often requiring extensive human annotation. Additionally, models’ reliance on real-world knowledge can lead to answer retrieval rather than logical deduction.
RealCQA-V2 offers a solution by building on human annotated structured visual data and constraining reasoning context to current visual input. To our knowledge this is the first work towards formalization of visual reasoning chains. By constraining ourselves to chart reasoning we attempt to build a benchmark that is ‘depth-first’. Instead of testing broad language based knowledge across multiple domains we focus on understanding a models capacity to model visual complexity seen in real world data. By design the charts require very similar initial logic to parse structure premises and then build on with data retrieval, and finally perform mathematical reasoning to correctly answer a question, we will discuss this in our methodology section 3.
@@ 2.3 Visual Reasoning over Charts
Charts convey complex information through visual elements like trend lines and legends, requiring interpretation of various visual tasks (see Figure 1). Charts have been studied either in a ‘dense’ setting, requiring localization-recognition of individual chart components and tabular data extraction or by downstream image to text tasks like Chart-QA, Chart summarization, Chart captioning etc.
@@
Dense Chart Prediction models like Cached [33] focus on chart structure prediction and proposed a custom local-global context modelling architecture based on Swin-Transformers. Chart-Ocr[34] and SpaDen [35] focuses on using keypoints for tabular data extraction, by recreating plot area components like bars, lines etc and end-to-end data extraction with legend mapping.
Chart Question Answering models like CRCT [36] use classification and regression predictors based on ‘Answer Type’. ChartQA[1] introduced template and human annotated question answers, baselined through offline tabular extraction using ChartOCR and answer prediction using VLT5. Matcha [37] utilizes the Pix2Struct model[9] to get an impressive jump in $20\%$ accuracy on ChartQA[1] by additional training on HTML structure and Math-QA.
Chart to Text models like Chartllama [38] use a generic visual encoder with LLAMA. ChartReader [39] further leverages a keypoint-based chart component detector with a T5-based text decoder. StructChart [40] showed novelty by reformulating chart data into triplets, effectively capturing both the structure and semantics inherent in the data. ChartAssistant [41] integrates Swin and ViT encoders with Bart and LLAMA language decoders, evaluating across multiple datasets. All of these chart to text works that report ‘reasoning’ accuracy are inherently ambiguous due to one-to-one evaluation of QA based visual reasoning.
Real World Charts still remain challenging. First observed in dense tasks [42] where synthetic (matplotlib) charts were easy to learn and models failed on real world (born digital pdf) charts. Recent study ChartXiv[43] has rehashed the same gap existing in the realm of LVLM’s as well. They undertook manual curation of around $\sim$1k charts from Arxiv publications and show the sota LLM’s which have a high performance (80%) on popular chart question answering benchmarks [1, 2] only achieve about 48% on their dataset. The RealCQA charts represent this visual complexity. Our proposed step by step text descriptors not only help in a nuanced evaluation for Real World visual reasoning but also helps improve the overall capacity for chart question answering. We will demonstrate this hypothesis in our results by doing a zero shot study on ChartXiv[43].
@@ 3 Chart FOL Premise Proving
We discuss our methodology for the proposed task. We want to build a system for evaluation of visual reasoning in charts from first principles as depicted in Figure 2. Our basic idea derives this theory from the way human beings read semantically arranged visual representations of data, i.e. charts. We start from the input an Image $I\in R^{3}$ and a question string. As a human we identify the valid components of the vision space in given context and then reason. By constraining ourselves to the domain of chart question answering, each logic sequence should capture how to read a chart(a fixed set of steps dependant on existing chart variables like title, label, tick , range etc), how to retrieve the data (can be generalized as keypoint estimation problem across chart types [34, 35]) and how to do the subsequent mathematical comparison(fixed set of predicates from QA). In the RealCQA dataset, each question context is confined to the template variables and predicates which were hand crafted, while the answer is either a chart attribute, data present in chart or mathematical reasoning over chart data. In our work we have systematically characterized these into a series of premise-conclusion pairs per question, visually grounded in the plot area. While ScienceQA and A-OKVQA had manual text annotators our dataset had manual vision annotators. Through multiple iterations of the [42] challenge (details in supplementary) our dataset consists of real world images painstakingly annotated for each component and data point. While the questions and premises remain template based however due to the nature of underlying scientific titles and symbols the result is a complex vocabulary.
@@ 3.1 VPP Task
We start by defining a Premise-Conclusion Pair, where each premise logically leads to a conclusion. While ScienceQA[28] and other multimodal COT datasets have multiple choice answers, we design these premises as binary FOL, where premises utilize quantifiers such as ‘for all’ ($\forall$) and ‘there exists’ ($\exists$), alongside a closed set of predicates, and the conclusions are True/False. Subsequently, we establish a Chain of Reasoning. This sequence involves linking each proposition to the next, ensuring that each step logically follows from the previous one as depicted as a flowchart in Figure LABEL:flow:chart-parse. The task culminates in Proving the Premises, where the final premise-conclusion, the original question—answer is derived from the rationales in the established reasoning chain. Successfully proving each premise confirms that the conclusion is a logical consequence of the rationales. The design choice of FOL premises enables us to also curate the graph representations of the logic sequences [44] using abstract syntax trees (AST) representing the canonical form for the sequence of rationales of each QA. Representing textual premises as trees where nodes correspond to variables or values and predicates, and edges represent logical or causal relationships between these entities enables problem formulation as a graph problem for visual logic verification studied independently of the natural language complexity. While we have released the AST representation of our dataset we currently focus on the NLP version for this work. In the future AST can be used for evaluating a visual premise generation (VPG) task. For now the VPP task aims to prove each premise as True/False. Figure 6 depicts a sample from the proposed dataset. We provide more details in supplementary.
@@ 3.2 Premise Conclusion Sequence Creation
We create four types of chart premises:
@@
    Structural Premises (SP) Identify the chart’s structure, categorized into 10 types (e.g., 0: Chart Type, 1: Y-Title, 2: X-Title, 3: Y-Range, 4: X-Range, 5: Categorical, 6: X-Tick, 7: Y-Tick, 8: # of Dataseries, 9: Legend Label).
@@
    Data Premises (DP) explicitly verify and retrieve data from the chart in visual space after plot area verification.
@@
    Reasoning Premises (RP) requiring logical deduction based on verified chart data and explicit calculations like comparing or transforming data values.
@@
While SP and DP are constructed per chart image, RP and MP are constructed for each question string. This creates unique combinations of logical sequences per question having similar initial logic and differing in the tail end. Figure 3 depicts SP by variable type, while every chart has chart type annotation, the chart text is annotated for a subset. This gives us a median of 4.5k SP in train per chart variable. Figure 4 depicts counts per chart image, $\sim$1.8k charts have legends and all 10 SP of 19k in train, while $\sim$5.6k charts have SP1-9. The charts with data annotations are a subset of the charts with text annotation. Figure 5 depicts our meaning of a ‘depth-first’ visual-reasoning dataset. In train we have $\sim$ 3.5k charts with a mean of 500 data premises and 101 reasoning premises per image. Around 300 charts have additional mean 149 MP’s per chart. Compared to any other existing multimodal dataset which has 4-15 text per image [28, 26]. This design enables a model to learn parsing the visual structure and data retrieval logic of a chart invariant of the reasoning logic. While the AST representation captures the canonical nature of the FOL, an FOL containing an expression such as ‘for all’ ($\forall$) and ‘there exists’ ($\exists$) leads to large combinatorial sets of DP/RP/MP for the same. This design truly challenges the visual capacity of a model. In forcing it to verify the same logic for every pertinent data point, we want to test the true visual parsing capability of LVLM’s. LLM’s are known to hallucinate when given the same string multiple times with different numerical values to compare. Here the model has to prove the same premise with different values grounded in visual space. Thus by law of large numbers, a model can not just find shortcuts through semantic language patterns and needs to correctly parse the visual content. Each premise template is used to generate a premise string and has true or false conclusion specific to query values of that chart used to generate the string. For valid conclusions we use actual variable value present in the groundtruth chart annotation. For invalid conclusions we randomly use any other relevant variable values present in the current chart (eg. out of range tick value, axis title instead of legend label etc) see example in supplementary. This setup effectively breaks down the original question into a comprehensible set of binary premise-conclusions, generated in a ratio of 1:3 true-to-false for each. This is done by design to bias the dataset with distractor (False) representations and prevent overfitting to actual (True) conclusion statements. In generating the final Premise strings we further add vocabulary complexity by choosing at random 1 of 3 paraphrased templates created using a pre-trained T5-Transformer from original handcrafted premise template. In ScienceQA they provide 3-4 MCQ answer choices (1 is correct) and in AOK-KVQA they provide 4 template answers (1 is correct) as distractors. Our choice of binary conclusions enable our evaluation metric.
Figure 3: SP by Chart Variable
Figure 4: SP Distribution (Hi-res in supplementary)
Figure 5: DP,RP,MP Distribution (Hi-res in supplementary)
Figure 6: Example data From RealCQA-V2
From Top-Left: A chart image, a question template(QT), its question string(QS), NLP Premises, FOL Premises, Atomic AST